Tom and David have 480 marbles between them. Tom arranges them in columns with 6 in each column, and David arranges them with 9 in each column. The total number of columns is 72. Express the above information in the form of an algebraic equation.
Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

6x+9(72-x)=480

Answer by stanbon(48558)

Tom and David have 480 marbles between them. Tom arranges them in columns with 6 in each column, and David arranges them with 9 in each column. The total number of columns is 72. Express the above information in the form of an algebraic equation.
---------------
Let T be the number of Tom's columns:
Then 6T is the number of marbles Tom has.
-------------
Let D be the number of Dave's columns:
Then 9D is the number of marbles Dave has.
----------------
6T+9D=480
D+T = 72
--------------
Solve the 2nd equation for D to get:
D = 72-T
-----------
Substitute into the 1st equation to solve for T
6T+9(72-T) = 480
6T+648-9T = 480
-3T = - 168
T = 56 (Tom has 56 columns)
D+56=72
D= 16 (Dave has 16 columns)
=====================
Cheers,
Stan H.
-